OpenStudy (anonymous):

Taye and Mike are raking leaves to earn some money. Taye can rake 2 lawns that are about the same size in 3 hours. Mike can rake 2 lawns that are about the same in 4 hours. How long would it take both boys to work together to rake 2 lawns? Write and solve an equation for this situation. Explain how to set up the equation, using w = rt.

OpenStudy (anonymous):

\[\frac{3\times 4}{3+4}\]

OpenStudy (anonymous):

ok. thats what i got but i did it differently, let me show u 2/3 (2) 4 /6 +1/2 (3) 3/6 =7/6 2/(7/6)= 12/7

OpenStudy (anonymous):

ok good, then no explanation needed if you it in the general case you can see that the answer is always in the same form \[\frac{xy}{x+y}\]

OpenStudy (anonymous):

yes. i see. thank you!